Comerpickles:

The volume of a sphere is 36π in³. What is the radius of the sphere?

dude:

We can use the volume formula \[V=\frac43πr^3\] We can work backwards and substitute the values we know

Comerpickles:

How'd you get 4/3? I mean, where did that come from?

dude:

That is the equation to find the volume of a sphere If you have a reference sheet it should be there

Comerpickles:

Oh okay so how do I identify the radius?

dude:

By substituting we get \[36\pi=\frac43πr^3\]

Comerpickles:

Ahh okay, let me do this real quick

Comerpickles:

Sorta like this? http://prntscr.com/i3x5wh

dude:

\(36\pi =\frac{4}{3}\pi r^3\) Multiply 3 on both sides \(108\pi =4\pi r^3\) Divide both sides by 4pi \(r^3=27\) We know that 27 is a perfect cube. So, \(r=3\)

Comerpickles:

So what do I do next? 36pi = 4/3 * 3^2?

dude:

Nothing else :P 3 is your radius

Comerpickles:

True, but they wanna know how many inches

Comerpickles:

So I just put in 3?

dude:

Yep, this is still in inches

Comerpickles:

Ohhh

Comerpickles:

How about this one? Just 3 more after this. http://prntscr.com/i3xcen

dude:

Heh, is it the same problem as last one except you were given a different value \(V=\frac43\pi 3^3\)

Comerpickles:

36pi

dude:

Right